OpenStudy (anonymous):

WILL MEDAL AND FAN !!!Solve for x. x - 3/2 =-5 (/ means division) A.-4 B.-7 C.-16 D.-13

OpenStudy (anonymous):

@King.Void.

OpenStudy (anonymous):

@Conqueror

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

So is this it? \(x - \dfrac{3}{2}{ = -5?}\)

OpenStudy (anonymous):

yes :0 thx that is much easier to read

OpenStudy (anonymous):

...-....

OpenStudy (anonymous):

Well, you would have to add \(\dfrac{3}{2}\) to both sides, so what is \(\dfrac{3}{2}\) + -5?

OpenStudy (anonymous):

13/2 =6.5 (6.5)

OpenStudy (anonymous):

Yes, so 6.5 is your answer. That's all you had to do. :)

OpenStudy (anonymous):

Wow :D thx here is your medal

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Oh, that isn't a choice... :/

OpenStudy (anonymous):

Oh, the answer was -3.5 but that isn't a choice either... :P

OpenStudy (anonymous):

divide 3/2 into both sides ?

OpenStudy (anonymous):

?? **###5